OpenStudy (anonymous):

PLEASE HELP!!!!!!!!! What is the 9th term of the geometric sequence 4, –20, 100, …?

OpenStudy (anonymous):

find the common ratio first\[a_n = a_1r^{n-1}\]\[a_2 = a_1r\]\[r = \frac{a_2}{a_1}\]\[r = \frac{-20}{4}\]\[r = -5\]after that find the 9th term\[a_9 = a_1r^8\]\[a_9 = 4\left( -5 \right)^8\]